A potentiometer wire of length 10 m and resistance 10 Ω per metre is connected in series with a resistance box and a 2 volt battery. If a potential difference of 100mV is balanced across the whole length of potentiometer wire, then the resistance introduced in the resistance will be
answer is 190ohms (im getting my answer as 1900 ohms pls help)

Grade:12

1 Answers

Sudarshan
5 years ago
       For potentiometer,        I= V÷R
                 I = 100 × 10^-3 ÷ 100
                 I = 10^-3
        Now,
                 I = EMF ÷ total resistance
          10^-3= 2 ÷( R + 100)
                 R=1900 Ohm